# Designing reaction-cross-diffusion systems with Turing and wave instabilities

### v1

@@ 1.1 Statement of the main results
For simplicity of presentation, we will focus on reaction-cross-diffusion systems on the real line, hence not being concerned with boundary conditions or geometry, as long as the eigenvalues of the Laplacian on the domain are sufficiently well-behaved. We note that this is essentially without loss of generality; see Murray2001 ; krause2021modern for discussion of this point in general, and (villar-sepulveda, , Remark 1) in particular. We do note that the use of Neumann or periodic boundary conditions can have important consequences for wave instabilities as these can give rise to different kinds of patterns, travelling waves or standing waves villar-sepulveda ; Villar-sepulveda-wave ; knobloch . However, these conditions do not impact the linear stability analysis. Given a fixed geometry and boundary conditions, one can simply scale the free matrix discussed below to fit a domain of a particular geometry, in any spatial dimension, assuming the Laplace (or more generally, the Laplace-Beltrami operator) spectrum is known and well-behaved for the given boundary conditions.
@@
where $\mathbf{u}\in\mathbb{R}^{n}$, $\mathbf{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, and $\mathbf{D}$ is the diffusion tensor of the system, which is not necessarily diagonal or linear. We will assume throughout that $n\geq 2$, though note that $n=2$ does not allow for wave instabilities even with cross-diffusion ritchie2022turing . We assume that sufficiently smooth and unique solutions exist, noting that the existence and regularity theory for these systems is much more intricate than for simpler reaction-diffusion models, with blowup and singularities having a significant literature choi2004existence ; le2005regularity ; seis-well ; choquet-well ; see lankeit2020facing for an introductory review to these complexities and their analysis. At a minimum, we will need $\mathbf{D}(\mathbf{u})$ to always have strictly positive real eigenvalues, and hence be invertible for all the values of its argument. Besides this assumption, we will henceforth neglect details regarding well-posedness and tacitly assume that everything at the nonlinear level is well-behaved.
We assume that Eq. 1 admits a spatially homogeneous steady state $\mathbf{P}$ such that $\mathbf{f}(\mathbf{P})=\mathbf{0}$. After linearizing about this steady state by writing $\mathbf{u}=\mathbf{P}+\varepsilon\mathbf{u}_{1}$ and keeping terms up to $\mathcal{O}(\varepsilon)$, we obtain
@@
Let $\mathbf{J}$ be a non-scalar stable matrix. Then, there exists a diagonalizable diffusion matrix $\mathbf{\hat{D}}$ such that Eq. 2 exhibits a Turing instability. Furthermore, if $n\geq 3$ and $\mathbf{J}$ has at least $1\leq\ell\leq 3$ distinct eigenvalues with at least $3-\ell$ corresponding generalized eigenvectors, then there is also a diagonalizable diffusion matrix $\mathbf{\hat{D}}$ such that Eq. 2 exhibits a wave instability.
@@
Figure 1 illustrates the key idea of the constructions for the case of a Turing instability. We construct a principal unstable submatrix that has the largest positive eigenvalue indicated by the green line. When $\delta=0$, a dispersion curve of the full system (shown in orange) tends asymptotically to this value as $\mu\to\infty$. We then increase $\delta$ to make the diffusion matrix nonsingular. This leads to the blue curves, which have the appropriate form of dispersion relation to ensure a Turing instability with a finite maximally unstable mode. The red curve illustrates the critical value of $\delta>0$ which gives this marginal stability criterion such that the dispersion curve is tangent with the horizontal axis. This represents the $\delta$-value at which a Turing bifurcation occurs. The blue dispersion curves thus represent a homotopy between the red and orange curves, parametrized by $\delta$. The proofs below will construct this homotopy explicitly, which will be illustrated in the examples.
@@ 2 Proof of Theorem 1
Let $\mathbf{\hat{D}}$ be diagonalizable matrix; specifically, there exists a non-singular matrix $\mathbf{Q}$ such that $\mathbb{D}:=\mathbf{Q}^{-1}\mathbf{\hat{D}}\mathbf{Q}$ is a diagonal matrix. Suppose further that $\mathbf{\hat{D}}$ has $\ell$ eigenvalues $\sigma_{i_{1}},\ldots,\sigma_{i_{\ell}}$ equal to zero, while all others are positive. Then, there exist $\ell$ eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ that tend to the eigenvalues of the principal submatrix of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ that comprises its $i_{1}^{\rm th},\ldots,i_{\ell}^{\rm th}$ rows and columns as $\mu\to\infty$. Meanwhile, the real parts of the eigenvalues associated with the principal submatrix of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ that excludes these rows and columns all tend to $\infty$- ∞ as $\mu\to\infty$.
Proof.
First, if $\ell=n$, then all the eigenvalues of $\mathbf{\hat{D}}$ are equal to zero. Therefore, as $\mathbf{\hat{D}}$ is diagonalizable, this implies it must be equal to the zero matrix, entailing that the eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ are equal to the eigenvalues of $\mathbf{J}$ for every $\mu>0$, which proves the result in this case.
@@
Let $\mathbf{w}_{m}=\mathbf{Q}^{-1}\mathbf{v}_{m}$. Therefore, (4) becomes
@@
which is equivalent to the linearization of the general system analyzed in villar-sepulveda , for $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ instead of $\mathbf{J}$. This implies that (villar-sepulveda, , Lemma 5) can be used to conclude the proof. ∎
To use the lemma above, we are going to focus on proving the existence of stable matrices $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ with unstable principal submatrices to conclude that the system can exhibit different diffusion-driven instabilities. Lemma 1 tells us that, to construct the diffusion matrices that allow the appearance of these instabilities, we need to follow two steps: first, we find the matrix $\mathbf{Q}$ of eigenvectors of $\mathbf{\hat{D}}$ such that $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ has an unstable principal submatrix; second, we change the eigenvalues of $\mathbf{\hat{D}}$ in order to control the convergence of the eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ as $\mu\to\infty$ (i.e. set the corresponding eigenvalues to asymptotically small values to induce the instability). While the case of zero diffusion coefficients corresponding to this principal unstable submatrix will lead to an ill-posed linear system, we can fix this by simply using small positive values instead of zeroes; see Fig. 1 for an illustration. An important point here is to understand what kind of structure we want to have for the unstable principle submatrix, and then use this structure to design the eigenvectors $\mathbf{Q}$ of the diffusion matrix to ensure that this submatrix appears within $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$.
@@ 2.1 Turing instabilities
Proof.
Let us denote the elements of $\mathbf{J}$ by
@@
This proves that $\mathbf{J}$ is similar to $\mathbf{K}$, which is a stable matrix with an unstable principal submatrix. Therefore, by choosing a suitable diagonal matrix $\mathbb{D}$, we can construct the matrix $\mathbf{\hat{D}}=\mathbf{Q}\mathbb{D}\mathbf{Q}^{-1}$ so that the result holds by using Lemma 1 (i.e. by choosing a sufficiently small bottom-right diagonal element to ensure a positive $\lambda$ for some $\mu>0$, thanks to the positive bottom-right element of $\mathbf{K}$).
Using Remark 1, and the fact that any non-scalar matrix $\mathbf{J}$ will have a $2\times 2$ non-scalar block in its Jordan Normal Form, we can immediately use Lemma 2 to construct a suitable matrix $\mathbf{\hat{D}}$ to induce Turing instabilities. The similarity matrix $\mathbf{Q}$ constructed in the proof of the Lemma is the key to proving the general result. Note that $\mathbf{K}$ was chosen in order to fulfill the conditions provided in Lemma 1, but there are infinitely many other matrices that are similar to $\mathbf{J}$ and have an unstable principal submatrix.
@@ 2.2 Wave instabilities
Let $\mathbf{J}$ be a non-scalar stable matrix and $n=3$ or $n=4$. If $\mathbf{J}$ has $1\leq\ell\leq 3$ distinct eigenvalues, then there exists a diffusion matrix $\mathbf{\hat{D}}$ so that (2) exhibits a wave instability provided $\mathbf{J}$ has at least $3-\ell$ generalized eigenvectors.
Proof.
We first consider the case $n=3$. Let us denote the elements of $\mathbf{J}$ by
@@
$\ell_{1,3}$ $=-\frac{\gamma\left(3z^{2}+6z+2\right)+\alpha\left(z^{3}+3z^{2}-2\right)}{\alpha^{2}+\gamma^{2}}$
@@
which is different from zero provided that $\ell=1$, and $r_{3,1}^{2}+r_{4,1}^{2}\neq 0$. Note that the $2\times 2$ bottom-right submatrix of $\mathbf{K}$ has eigenvalues $-\alpha\pm\alpha i$, which have a positive real part, as $\alpha<0$. As before, we can therefore construct a matrix $\mathbf{\hat{D}}$ to conclude the proof using the result in Lemma 1.
@@
Proof of Theorem 1.
First, we consider how to induce Turing instabilities. Given that $\mathbf{J}$ is non-scalar, we know there must exist a non-scalar $2\times 2$ sub-block of its Jordan normal form which is also non-scalar. We can apply Lemma 2 to find a matrix $\mathbf{K}$ similar to this sub-block, and by construction, this $\mathbf{K}$ will have a principle unstable submatrix that induces a Turing instability. We then use the construction given in Remark 1 to embed this similarity inside of the matrix $\mathbf{M}$ in that remark, and then apply Lemma 1 on the matrix $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$, which now has a principal unstable submatrix by construction. This Lemma allows us to set one of the eigenvalues of the diffusion matrix (i.e. an element of $\mathbb{D}$) to zero in order to excite the corresponding unstable principal submatrix and induce a Turing instability in the full system. Note that we can take all other elements of $\mathbb{D}$ as any arbitrary positive numbers in order to use Lemma 1.
The proof for inducing wave instabilities proceeds in an analogous way as above, except that we will make use of Lemma 3 to construct a matrix $\mathbf{K}$ similar to $\mathbf{J}$, instead. By assumption, $\mathbf{J}$ has $1\leq\ell\leq 3$ distinct eigenvalues with at least $3-\ell$ corresponding generalised eigenvectors. Hence, either $\mathbf{J}$ has a $3\times 3$ sub-block of its Jordan normal form that exactly fulfills this criterion as well, or it has a $4\times 4$ sub-block with two repeated complex conjugate eigenvalues. In either of these cases, we can again embed the appropriate matrix $\mathbf{K}$ from the proof of Lemma 3 into the matrix $\mathbf{M}$ from above in order to construct a suitable similarity matrix $\mathbf{Q}$ such that $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ has a principal unstable submatrix with complex-conjugate unstable eigenvalues. Again, through an analogous use of Lemma 1, we can set the appropriate two elements of $\mathbb{D}$ to zero to induce a wave instability in the full system. This concludes the proof.
We remark that, in practice, one needs the zero elements of $\mathbb{D}$ to be asymptotically small rather than strictly zero to avoid ill-posedness of the linear problem, but the existence of the instability will persist for some range of ‘small’ diffusion constants. We note that in both the diagonal matrices $\mathbb{D}$ and in the similarity matrices $\mathbf{Q}$ constructed above, one has, in principle, freedom over elements not affecting the principle unstable submatrix to use more general constructions to gain additional flexibility over the elements we have not exploited here. We will illustrate some aspects of this in Appendix A through alternative lemmas one can use to construct these matrices, and in the examples in Section 4.
@@ 3.1 Turing instabilities
Proof.
As $\mathbf{\hat{D}}$ is a diagonalizable non-scalar matrix with positive real eigenvalues and the elements of its diagonalization $\mathbb{D}$ are sorted, then there must exist at least one integer $11\leq\ell\leq n-$ 1 such that $\sigma_{\ell}<\sigma_{\ell+1}$.
@@
and $b,c,q>0$ are real constants to be determined. By checking its trace and determinant, one can see that $\mathbf{S^{*}}$ is stable. By construction, we have that $n-2$ eigenvalues of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}-\mu\mathbb{D}$ have the form $-1-\sigma_{i}$, for a corresponding integer $1\leq i\leq n$ such that $i\neq\ell,\ell+1$. On the other hand, by considering the two eigenvalues we have singled out via $\mathbf{S^{*}}$, we see that the equations that determine them are uncoupled from all others and, hence, they form an $n=2$ linear system which is precisely the classical Turing system, with $\mathbf{S^{*}}$ playing the role of the $2\times 2$ Jacobian matrix. Thus, the largest of the other two eigenvalues will be positive for some value of $\mu>0$ if the system satisfies the classical 2-component Turing conditions (see, e.g., Murray2001 ):
@@
Hence, we can always pick values of $b,c,$ and $q$ to satisfy these, given any pair of distinct eigenvalues $\sigma_{\ell}<\sigma_{\ell+1}$. Therefore, if we set $\mathbf{J}$ using this construction, we will induce a Turing instability in the original system given by Eq. 2. ∎

### v2

@@ 1.1 Statement of the main results
For simplicity of presentation, we will focus on reaction-cross-diffusion systems on the real line, hence not being concerned with boundary conditions or geometry, as long as the eigenvalues of the Laplacian on the domain are sufficiently well-behaved. We note that this is essentially without loss of generality in terms of problems posed on Riemannian manifolds with Neumann or periodic boundary conditions; see Murray2001 ; krause2021modern for discussion of this point in general, and (villar-sepulveda, , Remark 1) in particular. We do note that the use of Neumann or periodic boundary conditions can have important consequences for wave instabilities as these can give rise to different kinds of patterns, travelling waves or standing waves villar-sepulveda ; Villar-sepulveda-wave ; knobloch , and that other boundary conditions can induce a variety of behaviours maini1997boundary ; krause2021isolating . Given a fixed geometry and boundary conditions, one can simply scale the free matrix discussed below to fit a domain of a particular geometry, in any spatial dimension, assuming the Laplace (or more generally, the Laplace-Beltrami operator) spectrum is known and well-behaved (i.e. consisting solely of point eigenvalues) for the given boundary conditions.
@@
where $\mathbf{u}\in\mathbb{R}^{n}$, $\mathbf{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, and $\mathbf{D}:\mathbb{R}^{n}\to\mathbb{R}^{n}\times\mathbb{R}^{n}$ is the diffusion tensor of the system, which is not necessarily diagonal. We will assume throughout that $n\geq 2$, though note that $n=2$ does not allow for wave instabilities even with cross-diffusion ritchie2022turing . We assume that sufficiently smooth and unique solutions exist, noting that the existence and regularity theory for these systems is much more intricate than for simpler reaction-diffusion models, with blowup and singularities having a significant literature choi2004existence ; le2005regularity ; seis-well ; choquet-well ; see lankeit2020facing for an introductory review to these complexities and their analysis. We will henceforth neglect details regarding well-posedness and tacitly assume that everything at the nonlinear level is well-behaved.
We assume that Eq. 1 admits a spatially homogeneous steady state $\mathbf{P}$ such that $\mathbf{f}(\mathbf{P})=\mathbf{0}$, and that $\mathbf{D}(\mathbf{P})$ has strictly positive real eigenvalues, and hence is invertible for all the values of its argument. After linearizing about this steady state by writing $\mathbf{u}=\mathbf{P}+\varepsilon\mathbf{u}_{1}$ and keeping terms up to $\mathcal{O}(\varepsilon)$, we obtain
@@
Let $\mathbf{J}$ be a non-scalar stable matrix. Then, there exists a diagonalizable diffusion matrix $\mathbf{\hat{D}}$ such that Eq. 2 exhibits a Turing instability. Furthermore, if $n\geq 3$ and $\mathbf{J}$ has at least $1\leq\ell\leq 3$ distinct eigenvalues with at least 33-$\ell$ corresponding generalized eigenvectors, then there is also a diagonalizable diffusion matrix $\mathbf{\hat{D}}$ such that Eq. 2 exhibits a wave instability.
@@
Fig. 1 illustrates the key idea of the constructions for the case of a Turing instability. We construct a principal unstable submatrix that has the largest positive eigenvalue indicated by the green line. When $\delta=0$, a dispersion curve of the full system (shown in orange) tends asymptotically to this value as $\mu\to\infty$. We then increase $\delta$ to make the diffusion matrix nonsingular. This leads to the blue curves, which have the appropriate form of dispersion relation to ensure a Turing instability with a finite maximally unstable mode. The red curve illustrates the critical value of $\delta>0$, which gives this marginal stability criterion such that the dispersion curve is tangent to the horizontal axis. This represents the $\delta$-value at which a Turing bifurcation occurs. The blue dispersion curves thus represent a homotopy between the red and orange curves, parametrized by $\delta$. The proofs below will construct this homotopy explicitly, which will be illustrated in the examples.
@@ 2 Proof of Theorem 1
Let $\mathbf{\hat{D}}$ be diagonalizable matrix; specifically, there exists a non-singular matrix $\mathbf{Q}$ such that $\mathbb{D}:=\mathbf{Q}^{-1}\mathbf{\hat{D}}\mathbf{Q}$ is a diagonal matrix. Suppose further that $\mathbf{\hat{D}}$ has $\ell$ eigenvalues $\sigma_{i_{1}},\ldots,\sigma_{i_{\ell}}$ equal to zero, while all others are positive. Then, there exist $\ell$ eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ that tend to the eigenvalues of the principal submatrix of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ that comprises its $i_{1}^{\rm th},\ldots,i_{\ell}^{\rm th}$ rows and columns as $\mu\to\infty$. Meanwhile, the real parts of the eigenvalues associated with the principal submatrix of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ that excludes these rows and columns all tend to -$\infty$ as $\mu\to\infty$.
Proof: First, if $\ell=n$, then all the eigenvalues of $\mathbf{\hat{D}}$ are equal to zero. Therefore, as $\mathbf{\hat{D}}$ is diagonalizable, this implies it must be equal to the zero matrix, entailing that the eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ are equal to the eigenvalues of $\mathbf{J}$ for every $\mu>0$, which proves the result in this case.
@@
Let $\mathbf{w}_{m}=\mathbf{Q}^{-1}\mathbf{v}_{m}$. Therefore, Eq. 4 becomes
@@
which is equivalent to the linearization of the general system analyzed in villar-sepulveda , for $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ instead of $\mathbf{J}$. This implies that (villar-sepulveda, , Lemma 5) can be used to conclude the proof. To use the lemma above, we are going to focus on proving the existence of stable matrices $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ with unstable principal submatrices to conclude that the system can exhibit different diffusion-driven instabilities. Lemma 1 tells us that, to construct the diffusion matrices that allow the appearance of these instabilities, we need to follow two steps: first, we find the matrix $\mathbf{Q}$ of eigenvectors of $\mathbf{\hat{D}}$ such that $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ has an unstable principal submatrix; second, we change the eigenvalues of $\mathbf{\hat{D}}$ in order to control the convergence of the eigenvalues of $\mathbf{J}-\mu\mathbf{\hat{D}}$ as $\mu\to\infty$ (i.e. set the corresponding eigenvalues to asymptotically small values to induce the instability). While the case of zero diffusion coefficients corresponding to this principal unstable submatrix will lead to an ill-posed linear system, we can fix this by simply using small positive values instead of zeroes; see Fig. 1 for an illustration. An important point here is to understand what kind of structure we want to have for the unstable principle submatrix, and then use this structure to design the eigenvectors $\mathbf{Q}$ of the diffusion matrix to ensure that this submatrix appears within $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$.
@@ 2.1 Turing instabilities
Proof: Let us denote the elements of $\mathbf{J}$ by
@@
This proves that $\mathbf{J}$ is similar to $\mathbf{K}$, which is a stable matrix with an unstable principal submatrix. Therefore, by choosing a suitable diagonal matrix $\mathbb{D}$, we can construct the matrix $\mathbf{\hat{D}}=\mathbf{Q}\mathbb{D}\mathbf{Q}^{-1}$ so that the result holds by using Lemma 1 (i.e. by choosing a sufficiently small bottom-right diagonal element to ensure a positive $\lambda$ for some $\mu>0$, thanks to the positive bottom-right element of $\mathbf{K}$). Using Remark 1, and the fact that any non-scalar matrix $\mathbf{J}$ will have a $2\times 2$ non-scalar block in its Jordan Normal Form, we can immediately use Lemma 2 to construct a suitable matrix $\mathbf{\hat{D}}$ to induce Turing instabilities. The similarity matrix $\mathbf{Q}$ constructed in the proof of the Lemma is the key to proving the general result. Note that $\mathbf{K}$ was chosen in order to fulfill the conditions provided in Lemma 1, but there are infinitely many other matrices that are similar to $\mathbf{J}$ and have an unstable principal submatrix.
@@ 2.2 Wave instabilities
Let $\mathbf{J}$ be a non-scalar stable matrix and $n=3$ or $n=4$. If $\mathbf{J}$ has $1\leq\ell\leq 3$ distinct eigenvalues, then there exists a diffusion matrix $\mathbf{\hat{D}}$ so that Eq. 2 exhibits a wave instability provided $\mathbf{J}$ has at least 33-$\ell$ generalized eigenvectors.
Proof: We first consider the case $n=3$. Let us denote the elements of $\mathbf{J}$ by
@@
$\ell_{1,3}$ $=-\frac{\gamma\left(3z^{2}+6z+2\right)+\alpha\left(z^{3}+3z^{2}-2\right)}{\alpha^{2}+\gamma^{2}}.$
@@
which is different from zero provided that $\ell=1$, and $r_{3,1}^{2}+r_{4,1}^{2}\neq 0$. Note that the $2\times 2$ bottom-right submatrix of $\mathbf{K}$ has eigenvalues $\alpha\pm\alpha i$, which have a positive real part, as $\alpha<0$. As before, we can therefore construct a matrix $\mathbf{\hat{D}}$ to conclude the proof using the result in Lemma 1.
@@
Proof: [Proof of Theorem 1] First, we consider how to induce Turing instabilities. Given that $\mathbf{J}$ is non-scalar, we know there must exist a non-scalar $2\times 2$ sub-block of its Jordan normal form which is also non-scalar. We can apply Lemma 2 to find a matrix $\mathbf{K}$ similar to this sub-block, and by construction, this $\mathbf{K}$ will have a principle unstable submatrix that induces a Turing instability. We then use the construction given in Remark 1 to embed this similarity inside of the matrix $\mathbf{M}$ in that remark, and then apply Lemma 1 on the matrix $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$, which now has a principal unstable submatrix by construction. This Lemma allows us to set one of the eigenvalues of the diffusion matrix (i.e. an element of $\mathbb{D}$) to zero in order to excite the corresponding unstable principal submatrix and induce a Turing instability in the full system. Note that we can take all other elements of $\mathbb{D}$ as any arbitrary positive numbers in order to use Lemma 1.
The proof for inducing wave instabilities proceeds in an analogous way as above, except that we will make use of Lemma 3 to construct a matrix $\mathbf{K}$ similar to $\mathbf{J}$, instead. By assumption, $\mathbf{J}$ has $1\leq\ell\leq 3$ distinct eigenvalues with at least 33-$\ell$ corresponding generalised eigenvectors. Hence, either $\mathbf{J}$ has a $3\times 3$ sub-block of its Jordan normal form that exactly fulfills this criterion as well, or it has a $4\times 4$ sub-block with two repeated complex conjugate eigenvalues. In either of these cases, we can again embed the appropriate matrix $\mathbf{K}$ from the proof of Lemma 3 into the matrix $\mathbf{M}$ from above in order to construct a suitable similarity matrix $\mathbf{Q}$ such that $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}$ has a principal unstable submatrix with complex-conjugate unstable eigenvalues. Again, through an analogous use of Lemma 1, we can set the appropriate two elements of $\mathbb{D}$ to zero to induce a wave instability in the full system. This concludes the proof.
We remark that, in practice, one needs the zero elements of $\mathbb{D}$ to be asymptotically small rather than strictly zero to avoid ill-posedness of the linear problem, but the existence of the instability will persist for some range of ‘small’ diffusion constants. We note that in both the diagonal matrices $\mathbb{D}$ and in the similarity matrices $\mathbf{Q}$ constructed above, one has, in principle, freedom over elements not affecting the principal unstable submatrix to use more general constructions to gain additional flexibility over the elements we have not exploited here. We will illustrate some aspects of this in Appendix A through alternative lemmas one can use to construct these matrices, and in the examples in Section 4.
@@ 3.1 Turing instabilities
Proof: As $\mathbf{\hat{D}}$ is a diagonalizable non-scalar matrix with positive real eigenvalues and the elements of its diagonalization $\mathbb{D}$ are sorted, then there must exist at least one integer $1\leq\ell\leq n-1$ such that $\sigma_{\ell}<\sigma_{\ell+1}$.
@@
and $b,c,q>0$ are real constants to be determined. By checking its trace and determinant, one can see that $\mathbf{S^{*}}$ is stable. By construction, we have that n2n-2 eigenvalues of $\mathbf{Q}^{-1}\mathbf{J}\mathbf{Q}-\mu\mathbb{D}$ have the form $1-\sigma_{i}$, for a corresponding integer $1\leq i\leq n$ such that $i\neq\ell,\ell+1$. On the other hand, by considering the two eigenvalues we have singled out via $\mathbf{S^{*}}$, we see that the equations that determine them are uncoupled from all others and, hence, they form an $n=2$ linear system which is precisely the classical Turing system, with $\mathbf{S^{*}}$ playing the role of the $2\times 2$ Jacobian matrix. Thus, the largest of the other two eigenvalues will be positive for some value of $\mu>0$ if the system satisfies the classical 2-component Turing conditions (see, e.g., Murray2001 ):
@@
Hence, we can always pick values of $b,c,$ and $q$ to satisfy these, given any pair of distinct eigenvalues $\sigma_{\ell}<\sigma_{\ell+1}$. Therefore, if we set $\mathbf{J}$ using this construction, we will induce a Turing instability in the original system given by Eq. 2.